# Towards Comprehensive Multimodal Perception: Introducing the Touch-Language-Vision Dataset

### v1

@@ 6 Conclusion
In this work, we construct the first touch-language-vision dataset, TLV, featuring sentence-level descriptions for multimodal alignment. To demonstrate the effectiveness of the TLV dataset, we extended OpenCLIP and proposed TLV-Link, an unsupervised lightweight training approach. Preliminary experiments validate that the TLV dataset facilitates better alignment between touch and language. The proposed method may apply to specific scenarios, but there is room for improvement in terms of performance, and further enhancements are needed. Additionally, we intend to extend the application of TLV to more tasks to fully exploit its potential.

### v3

@@ 6 Conclusion
In this work, we construct the first touch-language-vision dataset, TLV, featuring sentence-level descriptions for multimodal alignment. To demonstrate the effectiveness of the TLV dataset, we extended OpenCLIP and proposed STLV-Align, an unsupervised lightweight training approach. Preliminary experiments validate that the TLV dataset facilitates better alignment between touch and language. The proposed method may apply to specific scenarios, but there is room for improvement in terms of performance, and further enhancements are needed. Additionally, we intend to extend the application of TLV to more tasks to fully exploit its potential.